Schister Systems uses the following data in its Cost-Volume-Profit analyses: Total Sales $ 380,000 Variable expenses 228,000 Contribution margin 152,000 Fixed expenses 116,000 Net operating income $ 36,000 What is total contribution margin if sales volume increases by 20%? Multiple Choice $152,000 $43,200 $182,400 $28,800
Answer #1

Contribution margin ratio = Contribution margin/Sales = 152000/380000 = 40%

Total Sales = 380000*1.2 = 456000

Total Contribution margin = 456000*40% = 182400

So answer is c) $182400